if N= 2.37 represents a measurement, then we assume an accuracy of 2.37+or- 0.005, Express the accuracy assumption using an absolute value inequality.
Answer by stanbon(75887)

if N= 2.37 represents a measurement, then we assume an
accuracy of 2.37 +or- 0.005, Express the accuracy assumption using an absolute value inequality.
-------
2.37-0.005<= N <= 2.37+0.005
--------
Subtract 2.37 along the line to get:
-0.005 <= N-237 <= +0.005
-----------------------------
|N-2.37| <= 0.005
-----------
Cheers,
Stan H.
--------